OpenStudy (anonymous):

Graph 3x + 9y = -21 and check on your calculator. What is the x-intercept, the y-intercept, and the slope?

OpenStudy (katrinakaif):

First convert that given equation into a y = mx + b format Can you do that for me?

OpenStudy (anonymous):

21(3)+9 = Y

OpenStudy (anonymous):

Right? o.o

OpenStudy (katrinakaif):

Hmmph. No. Can you try again?

OpenStudy (anonymous):

:/ Damnit. I know b = y intercept and m = the slope. Thats all I know. ; - ; Sooo...Y= 3x+9?

OpenStudy (katrinakaif):

Hmm Your equation is: 3x + 9y = -21 Your first step is to bring the 3x onto the other side of the equal sign. How do we do that?

OpenStudy (anonymous):

Divide because the only thing the x and the 3 have in common is they are being multiplied?

OpenStudy (katrinakaif):

We would only do that if we needed to have x alone.

OpenStudy (anonymous):

Oh. :/ Then we subtract nine from nine and -21?

OpenStudy (katrinakaif):

Ah, can you follow me on a Twiddla session? This mini box would take a while to explain :/

OpenStudy (anonymous):

Twiddla?

OpenStudy (katrinakaif):

Its a free online whiteboard